SQ(1) SQ(1) NAME sq−squeezeasortedwordlistunsq−unsqueezeasortedwordlist SYNOPSIS sq<infile>outfile unsq<infile>outfile DESCRIPTION sqcompressesasortedlistofwords(adictionary).Forexample: sort/usr/dict/words|sq|compress>words.sq.Z willcompressdictbyaboutafactorof4. unsquncompresstheoutputofsq.Forexample: compress-d<words.sq.Z|unsq|sort-f-owords willuncompressadictionarycompressedwithsq. Thesqueezingisachievedbyeliminatingcommonprefixes,andreplacingthemwithasinglecharacterwhichencodesthenumberofcharacterssharedwiththeprecedingword.Theprefixsizeis encodedasasingleprintablecharacter:0-9represent0-9,A-Zrepresent10-35,anda-zrepresent 36-61. AUTHOR MikeWexler SEEALSO compress(1),sort(1). LOCAL 1